angellalena:

A car was purchased for $19655 and depreciates about 4.7% a year, what is the value of the car after 8 years?

youdontknowme:

\[0.047*8*19655=7390.28\] This shows how much the car depreciates, so you take the original value and subtract the depreciation. \[19655-7390.28=12264.72\] The new value of the car after 8 years is $12,264.72

Vocaloid:

@angellalena the formula for exponential decay is A = A_0 * (1-r)^t where r = the rate of decay as a decimal, t is the number of years, and A_0 is the initial amount plug the values from the formula in, and solve for A note: the solution from the previous poster does not consider compounding, and is probably not what they're looking for